I have a right triangle with a hypotenuse measuring 25 ft. I need to "find the approximate lengths of the legs". They are represented by x. That is all the information given. I know I can't use pythagorean theorem, due to not knowing either leg length. I have been trying to find other examples, but all that I've found has the length of at least one leg given.
Answer by rothauserc(4718)

since each leg is represented by x, you have an isosceles right triangle,
x^2 +x^2 = 25^2
2x^2 = 625
x^2 = 625 / 2
x = square root ( 625 / 2)
x = 17.68